# The Poisson Midpoint Method for Langevin Dynamics: Provably Efficient Discretization for Diffusion Models

### v2

@@ 1 Introduction
The task of sampling from a target distribution is central to Bayesian inference, generative modeling, differential privacy and theoretical computer science [48, 20, 17, 25]. Sampling algorithms, based on the discretization of a stochastic differential equation (SDE) called the Langevin Dynamics, are widely used. The straightforward time discretization (i.e., Euler Maruyama discretization) of Langevin dynamics, called Langevin Monte Carlo (LMC), is popular due to its simplicity. The convergence properties of LMC have been studied extensively in the literature under various conditions on the target distribution [8, 12, 11, 45, 15, 33, 7, 1, 6, 9, 16, 30, 53, 5]. LMC can suffer from slow convergence to the target distribution, and often requires a large number of steps with a very fine time discretization (i.e., small step-size), making it prohibitively expensive.
@@ 1.1 Prior Work
Euler Maruyama discretization of SDEs is known to be inefficient and many powerful numerical integration techniques have been studied extensively ( [23, 32, 3, 26]). However, higher order methods such as the Runge-Kutta method require the existence and boundedness of higher order derivatives of the drift. The Randomized Midpoint Method for LMC (RLMC) was introduced for strongly log-concave sampling [40] and was further explored in [50, 18]. It was shown that RLMC, under certain conditions, can sample with a larger step size for fewer steps compared to Euler Maruyama and yet obtain the same accuracy. RLMC is popular due to its simplicity, and ease of implementation and does not require higher order bounded derivatives of the drift function. However, the current theoretical results are restricted to the case of strongly log-concave sampling, whereas non-log-concave sampling is of immense practical interest.
@@ 1.2 Our Contributions
(1) We design the Poisson Midpoint Method which discretizes SDEs by approximating $K$-steps of Euler-Maruyama discretization with step-size $\frac{\alpha}{K}$ by just one step with step-size $\alpha$. We show a strong error bound between these two processes under general conditions in Theorem 1 (no assumption on mixing, smoothness etc). This is based on a Central Limit Theorem (CLT) based method in [10] to analyze stochastic approximations of LMC.
@@ 1.3 Notation:
$X_{0:T}$ denotes $(X_{t})_{0\leq t\leq T}$ and $X_{K(0:T)}$ denotes $(X_{tK})_{0\leq t\leq T}$. $\mathbf{I}$ denotes identity matrix over $\mathbb{R}^{k\times k}$ whenever $k$ is clear from context. For any vector $\mathbf{x}\in\mathbb{R}^{k}$, $\|\mathbf{x}\|$ denotes its Euclidean norm. For any $a,b\in\mathbb{Z}$ and $a>b$, we take the $\sum_{t=a}^{b}$ to be 00, and the product $\prod_{t=a}^{b}$ to be $1$. Underdamped Langevin Dynamics happens in $\mathbb{R}^{2d}$. Here, we take vectors named $X$ (along with subscripts and superscripts) as $X=\begin{bmatrix}U&V\end{bmatrix}^{\intercal}$ where $U,V\in\mathbb{R}^{d}$ also carry the same subscripts and superscripts (e.g: $\tilde{X}_{4}$ corresponds to $\tilde{U}_{4},\tilde{V}_{4}$). In this case $\mathbf{I}$ represents identity matrix in $\mathbb{R}^{2d\times 2d}$ and $\mathbf{I}_{d}$ denotes the identity matrix in $\mathbb{R}^{d\times d}$. For any random variable $X$, we let $\mathsf{Law}(X)$ denote its probability measure. By $\mathsf{TV}(\mu,\nu)$ and $\mathsf{KL}\left(\mu\boldsymbol{\bigr{|}\bigr{|}}\nu\right)$ we denote the total variation distance and KL divergence (respectively) between two probability measure $\mu,\nu$. $O,\Omega,\Theta$ are standard Kolmogorov complexity notations whereas $\tilde{O},\tilde{\Omega},\tilde{\Theta}$ are same as $O,\Omega,\Theta$ up to poly-logarithmic factors in the problem parameters such as $\frac{1}{\epsilon},\frac{1}{\alpha},K,T,d$.
@@ 2 Problem Setup
Given a random vector $X_{0}\in\mathbb{R}^{d}$, consider an iterative, discrete time process $(X_{t})_{t\in\mathbb{N}\cup\{0\}}$ over $\mathbb{R}^{d}$, with step-size $\alpha>0$ given by:
$X_{t+1}=A_{\alpha}X_{t}+G_{\alpha}b(X_{t},t\alpha)+\Gamma_{\alpha}Z_{t}$ (1)
Where $A_{\alpha},G_{\alpha},\Gamma_{\alpha}$ are $d\times d$ matrix valued functions of the step-size $\alpha$ and $(Z_{t})_{t\geq 0}\stackrel{{\scriptstyle\mathsf{i.i.d.}}}{{\sim}}\mathcal{N}(0,\mathbf{I}_{d})$. $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the drift. Call this process $\mathsf{S}(A,G,\Gamma,b,\alpha)$. We consider Overdamped Langevin Monte Carlo (OLMC), Underdamped Langevin Monte Carlo (ULMC) and DDPMs as key examples.
@@
Consider Overdamped Langevin Dynamics for some $F:\mathbb{R}^{d}\to\mathbb{R}$:
$d\bar{X}_{\tau}=-\nabla F(\bar{X}_{\tau})d\tau+\sqrt{2}dB_{\tau}$ (2)
Here $B_{\tau}$ is the standard Brownian motion in $\mathbb{R}^{d}$. Under mild conditions on $F$ and $\bar{X}_{0}$, $\mathsf{Law}(\bar{X}_{\tau})\stackrel{{\scriptstyle\tau\to\infty}}{{\to}}\pi^{*}$ where $\pi^{\star}(X)\propto\exp(-F(X))$ is the stationary distribution.
Picking $A_{\alpha}=\mathbf{I}$, $G_{\alpha}=\alpha\mathbf{I}$, $\Gamma_{\alpha}=\sqrt{2\alpha}\mathbf{I}$ and $b(\mathbf{x},t\alpha)=-\nabla F(\mathbf{x})$ in Equation (1) gives us Euler-Maruyama discretization of Overdamped Langevin Dynamics: $X_{t}$ in (1) approximates $\bar{X}_{\alpha t}$ [38, 35]. OLMC is the canonical algorithm for sampling and has been studied under assumptions such as log-concavity of $\pi^{\star}$ [8, 12, 11] or that $\pi^{\star}$ satisfies isoperimetric inequalities [45, 15, 33, 7, 1].
@@
$A_{h}:=\begin{bmatrix}\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}\\ 0&e^{-\gamma h}\mathbf{I}_{d}\end{bmatrix},\quad G_{h}:=\begin{bmatrix}\tfrac{1}{\gamma}(h-\tfrac{1}{\gamma}(1-e^{-\gamma h}))\mathbf{I}_{d}&0\\ \tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}&0\end{bmatrix}\quad b(X_{t},t\alpha):=\begin{bmatrix}-\nabla F(U_{t})\\ 0\end{bmatrix}$
@@
This choice of $A_{h},\Gamma_{h},G_{h},b(,)$ in Equation (1) gives the Euler-Maruyama discretization of Underdamped Langevin Dynamics (a.k.a. the Kinetic Langevin Dynamics) studied extensively in Physics[14]):
$d\bar{U}_{\tau}$ $=\bar{V}_{\tau}d\tau;\qquad d\bar{V}_{\tau}=-\gamma\bar{V}_{\tau}-\nabla F(\bar{U}_{\tau})+\sqrt{2\gamma}dB_{\tau}$ (3)
@@
See [43, 5, 4] for further details. Here $p_{\tau}$ is the probability density of $e^{-t}X^{*}+\sqrt{1-e^{-2t}}Z$ where $X^{*}$ is drawn from the target and $Z$ is drawn from $\mathcal{N}(0,\mathbf{I})$ independently. The drift $\nabla\log p_{\tau}$ is learned via neural networks for discrete time instants $\tau_{0},\dots,\tau_{n-1}$ (usually $n=1000$). In practice, the iterations are written in the form  [^2] The time convention in the DDPM literature is reverse: $X_{t-1}=a_{t}X_{t}+b_{t}\nabla\log p_{\tau_{n-1-t}}(X_{t})+\sigma_{t}Z_{t}$.: $X_{t+1}=a_{t}X_{t}+b_{t}\nabla\log p_{\tau_{t}}(X_{t})+\sigma_{t}Z_{t}$ where $a_{t},b_{t},\sigma_{t}$ are chosen for best performance. Aside from the original choice in [20], many others have been proposed ([2, 42]). Since $A_{\alpha},G_{\alpha},\Gamma_{\alpha}$ in Equation (1) are time independent, we provide a variant of the Poisson Midpoint Method to suit DDPMs in Section A.1, along with a few other optimizations.
@@ 2.1 Technical Notes
We first illustrate the Randomized Midpoint Method [40, 18, 50] by applying it to OLMC (to obtain RLMC) to motivate our method (the Poisson Midpoint Method) and explain why we expect a quadratic speed up shown in Section 3. Overdamped Langevin Dynamics (2) satisfies:
[…]
@@
[…]
Intuitively, we expect that reducing the bias leads to a quadratic speed-up. Let $Z,Z^{\prime}\sim\mathcal{N}(0,1)$ and independent. For $\epsilon$ small enough it is easy to show that, $\mathsf{KL}\left(\mathsf{Law}(Z+\epsilon)\boldsymbol{\bigr{|}\bigr{|}}\mathsf{Law}(Z)\right)=\Theta(\epsilon^{2})$ whereas $\mathsf{KL}\left(\mathsf{Law}(Z+\epsilon Z^{\prime})\boldsymbol{\bigr{|}\bigr{|}}\mathsf{Law}(Z)\right)=\Theta(\epsilon^{4})$. We hypothesize that $Z_{t}+\text{error in integral}$ is closer to $\mathcal{N}(0,\mathbf{I})$ when the error term has a small mean and a large variance (as in RLMC) than when it has a large mean but 00 variance (as in LMC). However, rigorous analysis of RLMC has only been done under assumptions like strong log-concavity of the target distribution. This is due to the fact that $Z_{t}$ is dependent on the error in the integral, disallowing the argument above.
@@ 2.2 The Poisson Midpoint Method
We introduce the Poisson Midpoint Method (PLMC) which approximates $K$ steps of $\mathsf{S}(A,G,\Gamma,b,\frac{\alpha}{K})$ with step-size $\frac{\alpha}{K}$ with one step of which has a step-size $\alpha$. We denote this by $\mathsf{PS}(A,G,\Gamma,b,\alpha,K)$ and let its iterates be denoted by $(\tilde{X}_{tK})_{t\geq 0}$ or $(X_{t}^{\mathsf{P}})_{t\geq 0}$. Suppose $H_{t,i}\in\{0,1\}$ be any binary sequence and $Z_{tK+i}$ be a sequence of i.i.d. $\mathcal{N}(0,\mathbf{I})$ for $t,i\in\mathbb{N}\cup\{0\},0\leq i\leq K-1$. Given $\tilde{X}_{tK}$, we define the interpolation:
$\hat{\tilde{X}}_{tK+i}:=A_{\frac{\alpha i}{K}}\tilde{X}_{tK}+G_{\frac{\alpha i}{K}}b(\tilde{X}_{tK},t\alpha)+\sum_{j=0}^{i-1}A_{\frac{\alpha(i-j-1)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}$ (7)
Note that this interpolation is cheap since every one of $\hat{\tilde{X}}_{tK+i}$ can be computed with just one evaluation of the function $b()$. We then define the refined iterates for $10\leq i\leq K-$ 1 for a given $t$ as:
@@
We call our method Poisson Midpoint method since in OPTION 1 the set of midpoints $\{\alpha t+\frac{\alpha i}{K}:H_{t,i}=1\}$ converges to a Poisson process over $[\alpha t,\alpha(t+1)]$ as $K\to\infty$.
@@
$M_{k}=A_{\tfrac{\alpha(i_{k}-i_{k-1})}{K}}M_{k-1}+\Gamma_{\tfrac{\alpha(i_{k}-i_{k-1})}{K}}W_{t,k}$
@@
Step 3: For $k=1,\dots,N$, compute $\hat{\tilde{X}}_{tK+i_{k}}:=A_{\frac{\alpha i_{k}}{K}}X^{\mathsf{P}}_{t}+G_{\frac{\alpha i_{k}}{K}}b(X^{\mathsf{P}}_{t},\alpha t)+M_{k}$
Step 4: $\mathsf{Corr}:=K\sum_{k=1}^{N}G_{\tfrac{(K-1-i_{k})\alpha}{K}}(b(\hat{\tilde{X}}_{tK+i},\frac{(tK+i)\alpha}{K})-b(X^{\mathsf{P}}_{t},\alpha t))$
Step 5: $X_{t+1}^{\mathsf{P}}=A_{\alpha}X_{t}^{\mathsf{P}}+G_{\alpha}b(X^{\mathsf{P}}_{t},\alpha t)+M_{N+1}+\mathsf{Corr}\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad~{}(9)$
That is, the algorithm first generates the random mid-points $H_{t,i}$, computes the interpolation $\hat{\tilde{X}}_{tK+i}$ only when $H_{t,i}=1$ and then computes $b(\hat{\tilde{X}}_{tK+i},\frac{(tK+i)\alpha}{K})$ for these points. These computations are then combined to compute $\tilde{X}_{(t+1)K}$. In most applications, it is computationally easy to generate Gaussian random vectors and perform vector operations such as summation. However, the evaluation of the drift function $b()$ is expensive. Therefore, in this work, we consider the number of evaluations of the drift function as the measure of computational complexity. The following proposition establishes that each iteration of PLMC requires 2 evaluations of $b()$ in expectation.

### v3

@@ 1 Introduction
The task of sampling from a target distribution is central to Bayesian inference, generative modeling, differential privacy and theoretical computer science [48, 20, 17, 25]. Sampling algorithms, based on the discretization of a stochastic differential equation (SDE) called the Langevin Dynamics, are widely used. The straight forward time discretization (i.e, Euler Maruyama discretization) of Langevin dynamics, called Langevin Monte Carlo (LMC), is popular due to its simplicity. The convergence properties of LMC have been studied extensively in the literature under various conditions on the target distribution [8, 12, 11, 45, 15, 33, 7, 1, 6, 9, 16, 30, 53, 5]. LMC can suffer from slow convergence to the target distribution, and often requires a large number of steps with a very fine time discretization (i.e, small step-size), making it prohibitively expensive.
@@ 1.1 Prior Work
Euler Maruyama discretization of SDEs is known to be inefficient and many powerful numerical integration techniques have been studied extensively ( [23, 32, 3, 26]). However, higher order methods such as the Runge-Kutta method require the existence and boundedness of higher order derivatives of the drift. The Randomized Midpoint Method for LMC (RLMC) was introduced for strongly log-concave sampling [40] and was further explored in [50, 18]. RLMC is popular due to its simplicity, and ease of implementation and does not require higher order bounded derivatives of the drift function. However, the current theoretical results are restricted to the case of strongly log-concave sampling, whereas non-log-concave sampling is of immense practical interest.
@@ 1.2 Our Contributions
(1) We design the Poisson Midpoint Method which discretizes SDEs by approximating $K$-steps of Euler-Maruyama discretization with step-size $\frac{\alpha}{K}$ with just one step with step-size $\alpha$. We show a strong error bound between these two processes under general conditions in Theorem 1 (no assumption on mixing, smoothness etc). This is based on a Central Limit Theorem (CLT) based method in [10] to analyze stochastic approximations LMC.
@@ 1.3 Notation:
$X_{0:T}$ denotes $(X_{t})_{0\leq t\leq T}$ and $X_{K(0:T)}$ denotes $(X_{tK})_{0\leq t\leq T}$. $\mathbf{I}$ denotes identity matrix over $\mathbb{R}^{k\times k}$ whenever $k$ is clear from context. For any vector $\mathbf{x}\in\mathbb{R}^{k}$, $\|\mathbf{x}\|$ denotes its Euclidean norm. For any $a,b\in\mathbb{Z}$ and $a>b$, we take the $\sum_{t=a}^{b}$ to be 0, and the product $\prod_{t=a}^{b}$ to be $1$. Underdamped Langevin Dynamics happens in $\mathbb{R}^{2d}$. Here, we take vectors named $X$ (along with subscripts and superscripts) as $X=\begin{bmatrix}U&V\end{bmatrix}^{\intercal}$ where $U,V\in\mathbb{R}^{d}$ also carry the same subscripts and superscripts (e.g: $\tilde{X}_{4}$ corresponds to $\tilde{U}_{4},\tilde{V}_{4}$). In this case $\mathbf{I}$ represents identity matrix in $\mathbb{R}^{2d\times 2d}$ and $\mathbf{I}_{d}$ denotes the identity matrix in $\mathbb{R}^{d\times d}$. For any random variable $X$, we let $\mathsf{Law}(X)$ denote its probability measure. By $\mathsf{TV}(\mu,\nu)$ and $\mathsf{KL}\left(\mu\boldsymbol{\bigr|\bigr|}\nu\right)$ we denote the total variation distance and KL divergence (respectively) between two probability measure $\mu,\nu$. $O,\Omega,\Theta$ are standard Kolmogorov complexity notations whereas $\tilde{O},\tilde{\Omega},\tilde{\Theta}$ are same as $O,\Omega,\Theta$ up to poly-logarithmic factors in the problem parameters such as $\frac{1}{\epsilon},\frac{1}{\alpha},K,T,d$.
@@ 2 Problem Setup
Consider an iterative, discrete time process $(X_{t})_{t\in\mathbb{N}\cup\{0\}}$ over $\mathbb{R}^{d}$, with step-size $\alpha>0$ given by:
$X_{t+1}=A_{\alpha}X_{t}+G_{\alpha}b(X_{t},t\alpha)+\Gamma_{\alpha}Z_{t}.$ (1)
Here $A_{\alpha},G_{\alpha},\Gamma_{\alpha}$ are $d\times d$ matrix valued functions of the step-size $\alpha$ and $(Z_{t})_{t\geq 0}\stackrel{{\scriptstyle\mathsf{i.i.d.}}}{{\sim}}\mathcal{N}(0,\mathbf{I}_{d})$. $b:\mathbb{R}^{d}\to\mathbb{R}^{d}$ is the drift. Call this process $\mathsf{S}(A,G,\Gamma,b,\alpha)$. We consider Overdamped Langevin Monte Carlo (OLMC), Underdamped Langevin Monte Carlo (ULMC) and DDPMs as key examples.
@@
Here, $A_{\alpha}=\mathbf{I}$, $G_{\alpha}=\alpha\mathbf{I}$, $\Gamma_{\alpha}=\sqrt{2\alpha}\mathbf{I}$ and $b(\mathbf{x},t\alpha)=-\nabla F(\mathbf{x})$ for some $F:\mathbb{R}^{d}\to\mathbb{R}$. This is the Euler-Maruyama discretization of Overdamped Langevin Dynamics:
$d\bar{X}_{\tau}=-\nabla F(\bar{X}_{\tau})d\tau+\sqrt{2}dB_{\tau}.$ (2)
Here $B_{\tau}$ is the standard Brownian motion in $\mathbb{R}^{d}$. Under mild conditions on $F$ and $\bar{X}_{0}$, $\mathsf{Law}(\bar{X}_{\tau})\stackrel{{\scriptstyle\tau\to\infty}}{{\to}}\pi^{*}$ where $\pi^{\star}(X)\propto\exp(-F(X))$ is the stationary distribution. $X_{t}$ in (1) approximates $\bar{X}_{\alpha t}$ [38, 35]. OLMC is the canonical algorithm for sampling and has been studied under assumptions such as log-concavity of $\pi^{\star}$ [8, 12, 11] or that $\pi^{\star}$ satisfies isoperimetric inequalities [45, 15, 33, 7, 1].
@@
$A_{h}:=\begin{bmatrix}\mathbf{I}_{d}&\tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}\\ 0&e^{-\gamma h}\mathbf{I}_{d}\end{bmatrix};\kern 5.0ptG_{h}:=\begin{bmatrix}\tfrac{1}{\gamma}(h-\tfrac{1}{\gamma}(1-e^{-\gamma h}))\mathbf{I}_{d}&0\\ \tfrac{1}{\gamma}(1-e^{-\gamma h})\mathbf{I}_{d}&0\end{bmatrix};\kern 5.0ptb(X_{t},t\alpha):=\begin{bmatrix}-\nabla F(U_{t})\\ 0\end{bmatrix}$;
@@
This is the Euler-Maruyama discretization of Underdamped Langevin Dynamics (a.k.a. the Kinetic Langevin Dynamics) studied extensively in Physics[14]):
$d\bar{U}_{\tau}$ $=\bar{V}_{\tau}d\tau;\qquad d\bar{V}_{\tau}=-\gamma\bar{V}_{\tau}-\nabla F(\bar{U}_{\tau})+\sqrt{2}dB_{\tau}.$ (3)
@@
See [43, 5, 4] for further details. The drift $\nabla\log p_{\tau}$ is learned via neural networks for discrete time instants $\tau_{0},\dots,\tau_{n-1}$ (usually $n=1000$). In practice, the iterations are written in the form  [^2] The time convention in the DDPM literature is reverse: $X_{t-1}=a_{t}X_{t}+b_{t}\nabla\log p_{\tau_{n-1-t}}(X_{t})+\sigma_{t}Z_{t}$.: $X_{t+1}=a_{t}X_{t}+b_{t}\nabla\log p_{\tau_{t}}(X_{t})+\sigma_{t}Z_{t}$ where $a_{t},b_{t},\sigma_{t}$ are chosen for best performance. Aside from the original choice in [20], many others have been proposed ([2, 42]). Since $A_{\alpha},G_{\alpha},\Gamma_{\alpha}$ in Equation (1) are time independent, we provide a variant of the Poisson Midpoint Method to suit DDPMs in Section A.1, along with a few other optimizations.
@@ 2.1 Technical Notes
We illustrate the Randomized Midpoint Method [40, 18, 50] by applying it to OLMC (to obtain RLMC) to motivate our method and explain why we expect a quadratic speed up. Overdamped Langevin Dynamics (2) satisfies:
$\bar{X}_{(t+1)\alpha}=\bar{X}_{t\alpha}-\int_{t\alpha}^{(t+1)\alpha}\nabla F(\bar{X}_{s})ds+\sqrt{2}(B_{(t+1)\alpha}-B_{t\alpha}).$ (6)
Taking $X_{t}$ as the approximation to $\bar{X}_{t\alpha}$, LMC approximates the integral $\int_{t\alpha}^{(t+1)\alpha}\nabla F(\bar{X}_{s})ds$ with $\alpha\nabla F(X_{t})$, giving a ‘biased’ estimator to the integral (conditioned on $X_{t}=\bar{X}_{t\alpha}$). This gives the LMC updates $X_{t+1}=X_{t}-\alpha\nabla F(X_{t})+\sqrt{2\alpha}Z_{t};\quad Z_{t}\sim\mathcal{N}(0,\mathbf{I})$. RLMC chooses a uniformly random point in the interval $t\alpha,(t+1)\alpha]$ instead of initial point $t\alpha$ as described below:
@@
Notice that $\sqrt{2u\alpha}Z_{t,1}+\sqrt{2(1-u)\alpha}Z_{t,2}|u,X_{t}\sim\mathcal{N}(0,\mathbf{I})$, and $\mathbb{E}[\alpha\nabla F(X_{t+u})|X_{t},Z_{t,1},Z_{t,2}]=\alpha\int_{0}^{1}F(X_{t+s})ds$ which is a better approximation of the integral than $\alpha\nabla F(X_{t})$. Therefore RLMC provides a nearly unbiased approximation to the updates in Equation (6).
Intuitively, we expect that reducing the bias leads to a quadratic speed-up. Let $Z,Z^{\prime}\sim\mathcal{N}(0,1)$ and independent. For $\epsilon$ small enough it is easy to show that, $\mathsf{KL}\left(\mathsf{Law}(Z+\epsilon)\boldsymbol{\bigr|\bigr|}\mathsf{Law}(Z)\right)=\Theta(\epsilon^{2})$ whereas $\mathsf{KL}\left(\mathsf{Law}(Z+\epsilon Z^{\prime})\boldsymbol{\bigr|\bigr|}\mathsf{Law}(Z)\right)=\Theta(\epsilon^{4})$. We hypothesize that $Z_{t}+\text{error in integral}$ is closer to $\mathcal{N}(0,\mathbf{I})$ when the error term has a small mean and a large variance (as in RLMC) than when it has a large mean but 0 variance (as in LMC). However, rigorous analysis of RLMC has only been done under assumptions like strong log-concavity of the target distribution. This is due to the fact that $Z_{t}$ is dependent on the error in the integral, disallowing the argument above.
@@ 2.2 The Poisson Midpoint Method
We approximate $K$ steps of $\mathsf{S}(A,G,\Gamma,b,\frac{\alpha}{K})$ with step-size $\frac{\alpha}{K}$ with one step of our algorithm $\mathsf{PS}(A,G,\Gamma,b,\alpha,K)$ which has a step-size $\alpha$. The iterates of $\mathsf{PS}(A,G,\Gamma,b,\alpha,K)$ are denoted by $(\tilde{X}_{tK})_{t\geq 0}$ or $(X_{t}^{\mathsf{P}})_{t\geq 0}$. Suppose $H_{t,i}\in\{0,1\}$ be any binary sequence and $Z_{tK+i}$ be a sequence of i.i.d. $\mathcal{N}(0,\mathbf{I})$ for $t,i\in\mathbb{N}\cup\{0\},0\leq i\leq K-1$. Given $\tilde{X}_{tK}$, we define the cheap interpolation:
$\hat{\tilde{X}}_{tK+i}:=A_{\frac{\alpha i}{K}}\tilde{X}_{tK}+G_{\frac{\alpha i}{K}}b(\tilde{X}_{tK},t\alpha)+\sum_{j=0}^{i-1}A_{\frac{\alpha(i-j-1)}{K}}\Gamma_{\tfrac{\alpha}{K}}Z_{tK+j}.$ (7)
We then define the refined iterates for $0\leq i\leq K-1$ for a given $t$ as:
@@
We call our method Poisson Midpoint method since in OPTION 1 the set of midpoints $\{\alpha t+\frac{\alpha i}{K}:H_{t,i}=1\}$ converges to a Poisson process over $\alpha t,\alpha(t+1)]$ as $K\to\infty$.
@@
$M_{k}=A_{\tfrac{\alpha(i_{k}-i_{k-1})}{K}}M_{k-1}+\Gamma_{\tfrac{\alpha(i_{k}-i_{k-1})}{K}}W_{t,k}.$
@@
Step 3: For $k=1,\dots,N$, compute $\hat{\tilde{X}}_{tK+i_{k}}:=A_{\frac{\alpha i_{k}}{K}}X^{\mathsf{P}}_{t}+G_{\frac{\alpha i_{k}}{K}}b(X^{\mathsf{P}}_{t},\alpha t)+M_{k}$.
Step 4: $\mathsf{Corr}:=K\sum_{k=1}^{N}A_{\tfrac{(K-1-i_{k})\alpha}{K}}G_{\tfrac{\alpha}{K}}(b(\hat{\tilde{X}}_{tK+i},\frac{(tK+i)\alpha}{K})-b(X^{\mathsf{P}}_{t},\alpha t))$.
Step 5: $X_{t+1}^{\mathsf{P}}=A_{\alpha}X_{t}^{\mathsf{P}}+G_{\alpha}b(X^{\mathsf{P}}_{t},\alpha t)+M_{N+1}+\mathsf{Corr}.\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad~(9)$